# PruningBench: A Comprehensive Benchmark of Structural Pruning

### v3

@@ Abstract
Structural pruning has emerged as a promising approach for producing more efficient models. Nevertheless, the community suffers from a lack of standardized benchmarks and metrics, leaving the progress in this area not fully comprehended. To fill this gap, we present the first comprehensive benchmark, termed PruningBench, for structural pruning. PruningBench showcases the following three characteristics: 1) PruningBench employs a unified and consistent framework for evaluating the effectiveness of diverse structural pruning techniques; 2) PruningBench systematically evaluates 16 existing pruning methods, encompassing a wide array of models (e.g., CNNs and ViTs) and tasks (e.g., classification and detection); 3) PruningBench provides easily implementable interfaces to facilitate the implementation of future pruning methods, and enables the subsequent researchers to incorporate their work into our leaderboards. We provide an online pruning platform http://pruning.vipazoo.cn for customizing pruning tasks and reproducing all results in this paper. Codes will be made publicly on https://github.com/HollyLee2000/PruningBench.
@@ 1 Introduction
Model compression is an essential pursuit in the domain of machine learning, motivated by the necessity to strike a balance between model accuracy and computational efficiency. Various approaches have been developed to create more efficient models, including pruning [1], quantization [2], decomposition [3], and knowledge distillation [4, 5]. Among the multitude of compression paradigms, pruning has proven itself to be remarkably effective and practical [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. The aim of network pruning is to eliminate redundant parameters of a network to produce sparse models and potentially speed up the inference. Mainstream pruning approaches can be categorized into structurual pruning and unstructurual pruning. Unstructural pruning needs the support of special hardware or software to zero partial weights without modifying the network structure, whereas structural pruning physically removes grouped parameters of the networks without the requirement of specific hardware or software, thus getting a wider domain of applications in practice.
@@
Table 1: Experimental settings in some representative structural pruning methods. “#Comp.” indicates the number of competitors under comparison. “Params” and “FLOPs” indicate whether parameter count and FLOPs are controlled when compared with alternative methods. “Regularizer” means whether a sparsity regularizer is employed.
@@
Pitfall 3: Comparisons without controlling variables. Current methods usually present the changes in parameters [10, 23, 41, 42, 43, 44, 45, 28, 29], FLOPs [6, 15, 7, 24, 33, 35, 46, 47, 38, 11], or both [12, 13, 18, 20, 9, 19, 48, 30, 39] after pruning, but neglect the consistency of these variables when comparing different methods. Since accuracy, model size, and computational load all differ significantly after pruning, such comparisons without controlling variables can be hard to comprehend and leave the state of the field confusing.
@@
(3) PruningBench is designed as an expandable package that standardizes experimental settings and eases the integration of new algorithmic implementations. PruningBench provides straightforward interfaces for implementing importance criteria methods and sparsity regularizers, facilitating the development, evaluation and integration of future pruning algorithms into the leaderboards (further details about the interfaces can be referred to in the SUPP). Furthermore, our online platform enables users to customize pruning tasks by selecting models, datasets, methods, and hyperparameters, facilitating the reproducibility of the results presented in the paper.
Reproducibility. Leaderboards and online pruning platform are now available on the website http://pruning.vipazoo.cn. Benchmarking more models is still in progress. The code will be made publicly available soon.
@@ 3 PruningBench Settings
With the proposed PruningBench framework, we make a comprehensive study on existing structural pruning methods. We provide two leaderboards for each model-dataset combination, one for sparsifying-stage methods and the other for pruning-stage methods. The experimental settings in our benchmark are summarized as follows. For more details, please refer to the SUPP.
Models and Datasets. The benchmark now has been conducted on visual classification and detection tasks. For visual classification, we carry out the pruning experiments on the widely used CIFAR100 [51] and ImageNet [52] datasets, with ResNet18, ResNet50 [54], VGG19 [55] and ViT-small [49]. For visual detection, evaluations are conducted with YOLOv8 [50] on the COCO dataset [53]. The ResNet models for CIFAR100 are sourced from [56], and VGG models are sourced from [57]. For all these CIFAR models, the pretrained models used for pruning are trained by ourselves (see SUPP for the training details). For ImageNet experiments, the ResNet models and pretrained weights are obtained from the torchvision library [58], while the ViT-small model and its pretrained weight are sourced from the timm library [59]. For COCO experiments, the implementation and pretrained weight of the YOLOv8 model are obtained from the ultralytics library [60].
@@
Performance Metrics. The performances of different pruning methods are evaluated at several predefined speedup ratios. On CIFAR100, the speedup ratios are defined as {$2$x, $4$x, $8$x}. For large datasets, COCO and ImageNet, where tasks become more complex and tolerable pruning decreases, we adopt the speedup ratios {$2$x, $3$x, $4$x}. At each speedup ratio, pruning methods are compared in metrics of accuracy, parameters, pruning time, and regularizing time if sparsity regularizers are used.
Pruning Schemes. Currently there exist two pruning schemes: local pruning and global pruning. Local pruning removes a consistent proportion of parameters for each group in the network [18, 20, 21, 22, 9, 24, 25, 27, 11]. However, as the importance and the redundancy of parameters across layers differ largely, a consistent pruning strategy is usually suboptimal. In contrast to local pruning, global pruning removes structures from all available structures of a network until a specific speedup ratio is reached [19, 15, 23, 13, 26, 15], without constraining the pruning ratio across different groups to be consistent. However, global pruning may prune the entire group at high speedup ratio, leaving the model functionality broken down. To address these issues, we propose protected global pruning in this benchmark , which preserves at least 10% of the parameters within each group with global pruning. Experiments demonstrate that protected global protection yields comparable results at low (e.g., $2$x) speedup ratio and significantly superior performance at high (e.g., $4$x and $8$x) speedup ratio. Experimental results and discussions are deferred to Section 4.
Hyperparameters. When evaluating pruning-stage methods (i.e., importance criteria), the sparsifying stage is skipped. All involved hyperparameters in the fine-tuning stage are fixed to be the same. For sparsifying-stage methods, however, evaluation becomes more complex. Sparsifying-stage methods still rely on importance criteria at the pruning stage. For CNN experiments, we employ MagnitudeL2 [21] and BNScale [26] when benchmark sparsifying-stage methods, which are proven to be stable and data-agnostic. However, for ViT experiments, we only use MagnitudeL2 due to the incompatibility of the ViT architecture with BNScale. Moreover, different sparsity regularizers have different hyperparameters, which are specific to each case and exhibit substantial differences across diverse model-dataset tasks. In our experiments, we carefully tune the hyperparameters individually for each sparsity regularizer. For more hyperparameters of sparsifying, pruning and finetuning stages, please refer to the SUPP.
@@ 4 Benchmark Results and Discussions
2x OBD-C [18] N/A 1 78.35 78.68 +0.33 16.45 M (69.39%) 7.559s N/A
Taylor [19] N/A 2 78.35 78.51 +0.16 16.65 M (70.24%) 3.740s N/A
FPGM [20] N/A 3 78.35 78.37 +0.02 15.37 M (64.84%) 0.163s N/A
MagnitudeL2 [21] N/A 4 78.35 78.32 -0.03 16.63 M (70.17%) 0.136s N/A
BNScale [26] N/A 5 78.35 78.30 -0.05 15.96 M (67.32%) 0.141s N/A
ThiNet [25] N/A 6 78.35 78.14 -0.21 15.19 M (64.06%) 33.619s N/A
Random [22] N/A 7 78.35 77.97 -0.38 11.78 M (49.70%) 0.104s N/A
CP [24] N/A 8 78.35 77.80 -0.55 7.15 M (30.15%) 2m51s N/A
MagnitudeL1 [21] N/A 9 78.35 77.62 -0.73 16.91 M (71.34%) 0.137s N/A
OBD-Hessian [15] N/A 10 78.35 77.26 -1.09 7.83 M (33.03%) 5m5s N/A
LAMP [23] N/A 11 78.35 76.26 -2.09 16.21 M (68.37%) 0.150s N/A
HRank [9] N/A 12 78.35 76.13 -2.22 6.47 M (27.29%) 34m32s N/A
MagnitudeL2 [21] GroupLASSO [61] 1 78.35 78.73 +0.38 16.51 M (69.66%) 0.136s 3m5s
BNScale [26] BNScale [26] 2 78.35 78.36 +0.01 15.97 M (67.37%) 0.141s 2m14s
MagnitudeL2 [21] GroupNorm [15] 3 78.35 78.30 -0.05 15.03 M (63.41%) 0.136s 3m7s
BNScale [26] GroupLASSO [61] 4 78.35 78.24 -0.11 15.86 M (66.90%) 0.141s 2m38s
MagnitudeL2 [21] GrowingReg [11] 5 78.35 77.99 -0.36 16.61 M (70.06%) 0.136s 3m1s
@@
MagnitudeL2 [21] N/A 2 78.35 77.98 -0.37 10.71 M (45.19%) 0.136s N/A
BNScale [26] N/A 3 78.35 77.90 -0.45 10.53 M (44.41%) 0.141s N/A
MagnitudeL1 [21] N/A 4 78.35 77.82 -0.53 11.10 M (46.81%) 0.137s N/A
Taylor [19] N/A 5 78.35 77.69 -0.66 5.47 M (23.09%) 3.740s N/A
OBD-C [18] N/A 6 78.35 77.51 -0.84 5.84 M (24.64%) 7.559s N/A
Random [22] N/A 7 78.35 77.41 -0.94 5.95 M (25.11%) 0.104s N/A
ThiNet [25] N/A 8 78.35 77.23 -1.12 4.72 M (19.91%) 33.619s N/A
CP [24] N/A 9 78.35 75.68 -2.67 2.65 M(11.18%) 2m51s N/A
LAMP [23] N/A 10 78.35 75.52 -2.83 5.93 M (25.03%) 0.150s N/A
OBD-Hessian [15] N/A 11 78.35 75.49 -2.86 3.26 M (13.75%) 5m5s N/A
HRank [9] N/A 12 78.35 73.76 -4.59 1.69 M(7.11%) 34m32s N/A
BNScale [26] BNScale [26] 1 78.35 78.16 -0.19 10.37 M (43.75%) 0.141s 2m14s
MagnitudeL2 [21] GroupLASSO [61] 2 78.35 78.01 -0.34 10.79 M (45.53%) 0.136s 3m5s
BNScale [26] GroupLASSO [61] 3 78.35 77.90 -0.45 10.76 M (45.38%) 0.141s 2m38s
MagnitudeL2 [21] GroupNorm [15] 4 78.35 77.88 -0.47 9.84 M (41.51%) 0.136s 3m7s
MagnitudeL2 [21] GrowingReg [11] 5 78.35 77.86 -0.49 10.77 M (45.43%) 0.136s 3m1s
8x MagnitudeL1 [21] N/A 1 78.35 76.99 -1.36 6.82 M (28.77%) 0.137s N/A
MagnitudeL2 [21] N/A 2 78.35 76.38 -1.97 6.89 M (29.05%) 0.136s N/A
Random [22] N/A 3 78.35 76.13 -2.22 2.98 M (12.57%) 0.104s N/A
@@
BNScale [26] N/A 5 78.35 75.81 -2.54 6.69 M (28.22%) 0.141s N/A
OBD-C [18] N/A 6 78.35 75.78 -2.57 2.35 M (9.92%) 7.559s N/A
Taylor [19] N/A 7 78.35 75.38 -2.97 1.98 M (8.34%) 3.740s N/A
ThiNet [25] N/A 8 78.35 75.29 -3.06 1.58 M (6.68%) 33.619s N/A
OBD-Hessian [15] N/A 9 78.35 74.49 -3.86 1.66 M (7.02%) 5m5s N/A
LAMP [23] N/A 10 78.35 73.48 -4.87 3.62 M (15.27%) 0.150s N/A
CP [24] N/A 11 78.35 72.39 -5.96 0.97 M (4.07%) 2m51s N/A
HRank [9] N/A 12 78.35 70.54 -7.81 0.64 M (2.69%) 34m32s N/A
MagnitudeL2 [21] GrowingReg [11] 1 78.35 76.39 -1.96 7.00 M (29.52%) 0.136s 3m1s
MagnitudeL2 [21] GroupLASSO [61] 2 78.35 76.27 -2.08 7.09 M (29.90%) 0.136s 3m5s
MagnitudeL2 [21] GroupNorm [15] 3 78.35 75.93 -2.42 7.18 M (30.28%) 0.136s 3m7s
BNScale [26] GroupLASSO [61] 4 78.35 75.60 -2.75 7.19 M (30.32%) 0.141s 2m38s
BNScale [26] BNScale [26] 5 78.35 75.47 -2.88 6.90 M (29.12%) 0.141s 2m14s

### v5

@@ Abstract
Structural pruning has emerged as a promising approach for producing more efficient models. Nevertheless, the community suffers from a lack of standardized benchmarks and metrics, leaving the progress in this area not fully comprehended. To fill this gap, we present the first comprehensive benchmark, termed PruningBench, for structural pruning. PruningBench showcases the following three characteristics: 1) PruningBench employs a unified and consistent framework for evaluating the effectiveness of diverse structural pruning techniques; 2) PruningBench systematically evaluates 16 existing pruning methods, encompassing a wide array of models (e.g., CNNs and ViTs) and tasks (e.g., classification and detection); 3) PruningBench provides easily implementable interfaces to facilitate the implementation of future pruning methods, and enables the subsequent researchers to incorporate their work into our leaderboards. We provide an online pruning platform http://pruning.vipazoo.cn for customizing pruning tasks and reproducing all results in this paper. Leaderboard results can be available on https://github.com/HollyLee2000/PruningBench.
keywords:
network compression, structural pruning, regularization,benchmark
@@ 1 Introduction
Model compression is an essential pursuit in the domain of machine learning, motivated by the necessity to strike a balance between model accuracy and computational efficiency. Various approaches have been developed to create more efficient models, including pruning [1], quantization [2], decomposition [3], and knowledge distillation [4, 5, 6]. Among the multitude of compression paradigms, pruning has proven itself to be remarkably effective and practical [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The aim of network pruning is to eliminate redundant parameters of a network to produce sparse models and potentially speed up the inference. Mainstream pruning approaches can be categorized into structurual pruning and unstructurual pruning. Unstructured pruning typically involves directly zeroing partial weights without modifying the network structure; whereas structured pruning methods, although some require specific hardware support, can physically remove grouped parameters from the network, they effectively compress the network size, thus getting a wider domain of applications in practice.
@@
Table 1: Experimental settings in some representative structural pruning methods. “#Comp.” indicates the number of pruning methods compared with the proposed method in the original paper. “Params” and “FLOPs” indicate whether parameter count and FLOPs are controlled when compared with alternative methods. “Regularizer” means whether a sparsity regularizer is employed.
@@
Pitfall 3: Comparisons without controlling variables. Current methods usually present the changes in parameters [11, 24, 42, 43, 44, 45, 29, 30], FLOPs [7, 16, 8, 25, 34, 36, 46, 47, 12], or both [13, 14, 19, 21, 10, 20, 48, 40] after pruning, but neglect the consistency of these variables when comparing different methods. In different methods, the usage of parameters and FLOPS is inconsistent, which makes it impossible to compare the pruning results. Since accuracy, model size, and computational load all differ significantly after pruning, such comparisons without controlling variables can be hard to comprehend and leave the state of the field confusing.
@@
(3) PruningBench is designed as an expandable package that standardizes experimental settings and eases the integration of new algorithmic implementations. PruningBench provides straightforward interfaces for implementing importance criteria methods and sparsity regularizers, facilitating the development, evaluation and integration of future pruning algorithms into the leaderboards (further details about the interfaces can be referred to in the A.2.1). Furthermore, our online platform enables users to customize pruning tasks by selecting models, datasets, methods, and hyperparameters, facilitating the reproducibility of the results presented in the paper.
Reproducibility. Leaderboards and online pruning platform are now available on the website http://pruning.vipazoo.cn. Benchmarking more models is still in progress.
2 Related work
PruningBench categorizes current structural pruning methods into importance criteria and sparsity regularizers. Importance criteria assess the importance of filters within a neural network, identifying redundant filters or channels which need to be pruned, whereas sparsity regularizers aim to learn structured sparse networks by imposing sparse constraints on loss functions and zeroing out certain weights during training.
The sparsity regularizers can be applied to Batch Normalization (BN) parameters [54] if the networks contain batch normalization layers [27, 38, 39, 31], and then the BN parameters are used to indicate the pruning decision of structures such as channels or filters. Sparsity regularizers can also be directly applied to filters [54, 12, 16, 55, 28]. Group Lasso regularization [55, 28] is commonly used to sparsify filters in a structured manner. Growing Regularization (GREG) [12] exploits regularization under a growing penalty and uses two algorithms. More recently, GroupNorm [16] promotes sparsity across all grouped layers, convering convolutions, batch normalizations and fully-connected layers.
Importance criteria can be divided into two approaches: data-free methods and data-driven methods. Data-free methods rely solely on the existing weight information of the network and do not depend on input data, making their pruning results deterministic. These methods can be classified into four categories: weight-norm, weight-correlation, BN-based, and random. Weight-norm methods [22, 24, 56] prune based on the norms of weight values. Representative works, such as MagnitudeL1 [22], MagnitudeL2 [22], and LAMP [24], consider filters with smaller norms to have weak activation, thus contributing less to the final classification decision [54]. Weight-correlation methods [21, 33, 34, 48, 57] prune based on the relationships between weight values. For instance, FPGM identifies filters close to the geometric median to be redundant, as they represent common information shared by all filters in the same layer and should be removed. BN-based methods [12, 16, 55, 28] prune based on the weights of BN layers. BNScale [27] directly uses the scaling parameter $\gamma$ of the BN layer to compute the importance scores, while Kang et al. [40] also consider shifting parameters $\beta$. Random methods [23] perform pruning in a random manner.
In contrast, data-driven methods are pruning techniques that require input samples, making their results non-deterministic and dependent on the quality of the input data. These methods can be categorized into activation-based and gradient-based approaches. Activation-based methods [25, 58, 10, 26, 59, 35, 29, 30] utilize activation maps for pruning. For example, CP [25] and HRank [10] evaluate channel importance of current layer using reconstruction error and activation map decomposition, respectively. ThiNet [26], on the other hand, uses activation maps from the next layer to guide the pruning of the current layer. Gradient-based methods [60, 61, 19, 16, 20, 62, 63] rely on gradients or Hessian information to perform pruning. Methods that rely solely on gradients, such as Taylor-FO [20] and Mol-16 [37], can obtain importance scores from backpropagation without requiring additional memory. Conversely, hessian-based methods, such as OBD-Hessian [16] and OBD-C [19], require calculating second-order derivatives, which are computationally prohibitive.
@@ 4 PruningBench Settings
With the proposed PruningBench framework, we make a comprehensive study on existing structural pruning methods. We provide two leaderboards for each model-dataset combination, one for sparsifying-stage methods and the other for pruning-stage methods. The experimental settings in our benchmark are summarized as follows.
Models and Datasets. The benchmark now has been conducted on visual classification and detection tasks. For visual classification, we carry out the pruning experiments on the widely used CIFAR100 [51] and ImageNet [52] datasets, with ResNet18, ResNet50 [64], VGG19 [65] and ViT-small [49]. For visual detection, evaluations are conducted with YOLOv8 [50] on the COCO dataset [53]. The ResNet models for CIFAR100 are sourced from [66], and VGG models are sourced from [67]. For all these CIFAR models, the pretrained models used for pruning are trained by ourselves (see A.1.2 for the training details). For ImageNet experiments, the ResNet models and pretrained weights are obtained from the torchvision library [68], while the ViT-small model and its pretrained weight are sourced from the timm library [69]. For COCO experiments, the implementation and pretrained weight of the YOLOv8 model are obtained from the ultralytics library [70].
@@
Performance Metrics. The performances of different pruning methods are evaluated at several predefined speedup ratios.On CIFAR100, the speedup ratios are defined as {$2$x, $4$x, $8$x}. For large datasets, COCO and ImageNet, where tasks become more complex and tolerable pruning decreases, we adopt the speedup ratios {$2$x, $3$x, $4$x}. Note that speedup ratio represents the difference in FLOPS between the pruned model and the original model. At each speedup ratio, pruning methods are compared in metrics of accuracy, parameters, pruning time, and regularizing time if sparsity regularizers are used.
Pruning Schemes. Currently there exist two pruning schemes: local pruning and global pruning. Local pruning removes a consistent proportion of parameters for each group in the network [19, 21, 22, 23, 10, 25, 26, 28, 12]. However, as the importance and the redundancy of parameters across layers differ largely, a consistent pruning strategy is usually suboptimal. In contrast to local pruning, global pruning removes structures from all available structures of a network until a specific speedup ratio is reached [20, 16, 24, 14, 27, 16], without constraining the pruning ratio across different groups to be consistent. However, global pruning may prune the entire group at a high speedup ratio, leaving the model functionality broken down. To address these issues, we propose protected global pruning in this benchmark , which preserves at least 10% of the parameters within each group with global pruning. Experiments demonstrate that protected global protection yields comparable results at low (e.g., $2$x) speedup ratio and significantly superior performance at high (e.g., $4$x and $8$x) speedup ratio. Like other studies, we also adopt the pruning strategy that controls FLOPS. Experimental results and discussions are deferred to Section 5.
Hyperparameters. When evaluating pruning-stage methods (i.e., importance criteria), the sparsifying stage is skipped. All involved hyperparameters in the fine-tuning stage are fixed to be the same. For sparsifying-stage methods, however, evaluation becomes more complex. Sparsifying-stage methods still rely on importance criteria at the pruning stage. For CNN experiments, we employ MagnitudeL2 [22] and BNScale [27] when benchmarking sparsifying-stage methods, which are proven to be stable and data-agnostic. However, for ViT experiments, we only use MagnitudeL2 due to the incompatibility of the ViT architecture with BNScale. Moreover, different sparsity regularizers have different hyperparameters, which are specific to each case and exhibit substantial differences across diverse model-dataset tasks. In our experiments, we carefully tune the hyperparameters individually for each sparsity regularizer. For more hyperparameters of sparsifying, pruning and finetuning stages, please refer to the A.1.3 A.1.4.
@@ 5 Benchmark Results and Discussions
2x OBD-C N/A 1 78.35 78.68 +0.33 16.45 M (69.39%) 7.559s N/A
Taylor N/A 2 78.35 78.51 +0.16 16.65 M (70.24%) 3.740s N/A
FPGM N/A 3 78.35 78.37 +0.02 15.37 M (64.84%) 0.163s N/A
MagnitudeL2 N/A 4 78.35 78.32 -0.03 16.63 M (70.17%) 0.136s N/A
BNScale N/A 5 78.35 78.30 -0.05 15.96 M (67.32%) 0.141s N/A
ThiNet N/A 6 78.35 78.14 -0.21 15.19 M (64.06%) 33.619s N/A
Random N/A 7 78.35 77.97 -0.38 11.78 M (49.70%) 0.104s N/A
CP N/A 8 78.35 77.80 -0.55 7.15 M (30.15%) 2m51s N/A
MagnitudeL1 N/A 9 78.35 77.62 -0.73 16.91 M (71.34%) 0.137s N/A
OBD-Hessian N/A 10 78.35 77.26 -1.09 7.83 M (33.03%) 5m5s N/A
LAMP N/A 11 78.35 76.26 -2.09 16.21 M (68.37%) 0.150s N/A
HRank N/A 12 78.35 76.13 -2.22 6.47 M (27.29%) 34m32s N/A
MagnitudeL2 GroupLASSO 1 78.35 78.73 +0.38 16.51 M (69.66%) 0.136s 3m5s
BNScale BNScale 2 78.35 78.36 +0.01 15.97 M (67.37%) 0.141s 2m14s
MagnitudeL2 GroupNorm 3 78.35 78.30 -0.05 15.03 M (63.41%) 0.136s 3m7s
BNScale GroupLASSO 4 78.35 78.24 -0.11 15.86 M (66.90%) 0.141s 2m38s
MagnitudeL2 GrowingReg 5 78.35 77.99 -0.36 16.61 M (70.06%) 0.136s 3m1s
@@
MagnitudeL2 N/A 2 78.35 77.98 -0.37 10.71 M (45.19%) 0.136s N/A
BNScale N/A 3 78.35 77.90 -0.45 10.53 M (44.41%) 0.141s N/A
MagnitudeL1 N/A 4 78.35 77.82 -0.53 11.10 M (46.81%) 0.137s N/A
Taylor N/A 5 78.35 77.69 -0.66 5.47 M (23.09%) 3.740s N/A
OBD-C N/A 6 78.35 77.51 -0.84 5.84 M (24.64%) 7.559s N/A
Random N/A 7 78.35 77.41 -0.94 5.95 M (25.11%) 0.104s N/A
ThiNet N/A 8 78.35 77.23 -1.12 4.72 M (19.91%) 33.619s N/A
CP N/A 9 78.35 75.68 -2.67 2.65 M(11.18%) 2m51s N/A
LAMP N/A 10 78.35 75.52 -2.83 5.93 M (25.03%) 0.150s N/A
OBD-Hessian N/A 11 78.35 75.49 -2.86 3.26 M (13.75%) 5m5s N/A
HRank N/A 12 78.35 73.76 -4.59 1.69 M(7.11%) 34m32s N/A
BNScale BNScale 1 78.35 78.16 -0.19 10.37 M (43.75%) 0.141s 2m14s
MagnitudeL2 GroupLASSO 2 78.35 78.01 -0.34 10.79 M (45.53%) 0.136s 3m5s
BNScale GroupLASSO 3 78.35 77.90 -0.45 10.76 M (45.38%) 0.141s 2m38s
MagnitudeL2 GroupNorm 4 78.35 77.88 -0.47 9.84 M (41.51%) 0.136s 3m7s
MagnitudeL2 GrowingReg 5 78.35 77.86 -0.49 10.77 M (45.43%) 0.136s 3m1s
8x MagnitudeL1 N/A 1 78.35 76.99 -1.36 6.82 M (28.77%) 0.137s N/A
MagnitudeL2 N/A 2 78.35 76.38 -1.97 6.89 M (29.05%) 0.136s N/A
Random N/A 3 78.35 76.13 -2.22 2.98 M (12.57%) 0.104s N/A
@@
BNScale N/A 5 78.35 75.81 -2.54 6.69 M (28.22%) 0.141s N/A
OBD-C N/A 6 78.35 75.78 -2.57 2.35 M (9.92%) 7.559s N/A
Taylor N/A 7 78.35 75.38 -2.97 1.98 M (8.34%) 3.740s N/A
ThiNet N/A 8 78.35 75.29 -3.06 1.58 M (6.68%) 33.619s N/A
OBD-Hessian N/A 9 78.35 74.49 -3.86 1.66 M (7.02%) 5m5s N/A
LAMP N/A 10 78.35 73.48 -4.87 3.62 M (15.27%) 0.150s N/A
CP N/A 11 78.35 72.39 -5.96 0.97 M (4.07%) 2m51s N/A
HRank N/A 12 78.35 70.54 -7.81 0.64 M (2.69%) 34m32s N/A
MagnitudeL2 GrowingReg 1 78.35 76.39 -1.96 7.00 M (29.52%) 0.136s 3m1s
MagnitudeL2 GroupLASSO 2 78.35 76.27 -2.08 7.09 M (29.90%) 0.136s 3m5s
MagnitudeL2 GroupNorm 3 78.35 75.93 -2.42 7.18 M (30.28%) 0.136s 3m7s
BNScale GroupLASSO 4 78.35 75.60 -2.75 7.19 M (30.32%) 0.141s 2m38s
BNScale BNScale 5 78.35 75.47 -2.88 6.90 M (29.12%) 0.141s 2m14s